1.) Two objects have masses of 300 kg and 500 kg. Their centers are separated by distance of 25 m. Find the gravitational attraction between them.


1
Expert's answer
2021-12-16T11:31:21-0500

Given:

"m_1=300\:\rm kg"

"m_2=500\:\rm kg"

"r=25\:\rm m"

"G=6.67*10^{-11}\:\rm N\cdot m^2/kg^2"


The universal law of gravity says

"F=G\frac{m_1m_2}{r^2}"

"F=6.67*10^{-11}\frac{300*500}{25^2}=1.6*10^{-8}\:\rm N"